OpenStudy (deadshot):

Solve. 1 over 25 = 5^(x + 4)?

OpenStudy (deadshot):

\[\frac{ 1 }{ 25 }=5^{x+4}\]

OpenStudy (campbell_st):

same process as the last question the base on both sides of the equation is 5 rewriting \[5^2 = 5^{x + 4}\] equate the powers and solve for x

OpenStudy (amistre64):

-2, but yes

OpenStudy (campbell_st):

so is it 1/25 or 25...?

OpenStudy (amistre64):

if you couldnt determine a like base, logging it would suffice

OpenStudy (campbell_st):

oops it should be misread the question. \[5^{-2} = 5^{x + 4}\]

OpenStudy (amistre64):

ln(1/25) = ln(5^(x+4)) ln(1} - ln(25) = ln(5) (x+4) - ln(25)/ln(5) = x+4 - ln(25)/ln(5) - 4 = x

OpenStudy (campbell_st):

why not just equate the powers since the bases are the same. -2 = x + 4

OpenStudy (amistre64):

you can, but spose you didnt notice a similarity of bases ...

OpenStudy (deadshot):

so x=-6, right?

OpenStudy (campbell_st):

looks like it...

OpenStudy (deadshot):

Okay, thanks!